Your firm needs a computerized machine tool lathe which costs $42,000 and requires $11,200 in maintenance for each year of its 4-year life. After three years, this machine will be replaced. The machine falls into the MACRS 5-year class life category. Assume a tax rate of 25 percent and a discount rate of 12 percent. If the lathe can be sold for $4,200 at the end of year 4, what is the after-tax salvage value?

Is there a way to solve in Excel?
Answer #1

rate positively ..

Computation of after tax salvage value -
a Book value after 4 year =     7,257.60
42000*(11.52%+5.76%)
b Sales price =     4,200.00
c=b-a Profit on sale = (3,057.60)
d=c*25% Tax on profit =      (764.40)
e=b-d Post tax salvage value =     4,964.40
Ans =     4,964.40